A car travels 200 mi. A second car, traveling 10mph faster than the first car, makes the same trip in one hour less time. Find the speed of each car.
Answer by TimothyLamb(4379)

D = 200
s = d / t
d = s * t
---
car a:
T = time for car a
200 = sT
---
car b:
200 = (s + 10)(T - 1)
---
sT = (s + 10)(T - 1)
sT = sT - s + 10T - 10
s = 10T - 10
---
200 = sT
200 = (10T - 10)T
10TT - 10T - 200 = 0
10T^2 - 10T - 200 = 0
---
the above quadratic equation is in standard form, with a=10, b=-10, and c=-200
---
to solve the quadratic equation, by using the quadratic formula, plug this:
10 -10 -200
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
the two real roots (i.e. the two solutions), of the quadratic are:
T = 5
T = -4
---
the negative root doesn't make sense for time, so use the positive root:
---
T = 5
s = 10T - 10
s = 10(5) - 10
s = 40
---
answer:
speed of car a: 40 mph
speed of car b: 50 mph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php